OpenStudy (anonymous):

x(2x+5)=0 .. Zero property plllsz.

OpenStudy (anonymous):

jjkm18 have done a few of these already... what is your first step?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

If two numbers multiply together to get 0, then one of them is 0.. Do you understand that?

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

Tell me two numbers that multiply together to get 0

OpenStudy (anonymous):

2x0 & 3x0.

OpenStudy (anonymous):

OK, do you agree that one of them had to be 0

OpenStudy (anonymous):

OH, Now I get it ..

OpenStudy (anonymous):

So it's zero?

OpenStudy (anonymous):

OK, so you have two numbers that are multiplied together to get 0 x and 2x + 5 so either x = 0 or the 2x + 5 = 0 x = 0 is already solved 2x + 5 = 0 (you must solve this for x) 2x = -5 x = -5/2 SO x = 0 or x = -5/2

OpenStudy (anonymous):

Dude, how do yu even understand math that much.. It's sooo confusinq.

OpenStudy (anonymous):

I'm a teacher LOL that's my job.

OpenStudy (anonymous):

Ohhhh, Hahaha. Math is sooo hard.

OpenStudy (anonymous):

Once you get it, it isn't... It can be very frustrating each time you learn a new topic.

OpenStudy (anonymous):

Hahaha. I have to get taught like 20 times to get it, Lol.